a) If two charges +7 x10-9C and -6 x10-9C are separated 0.04m suspended in air, at which point on the line that join both charges the magnitude of the electric field will be E = 0 N/C?



1
Expert's answer
2021-10-08T09:23:20-0400
"E_1+E_2=0,""E_1=-E_2,""\dfrac{kq_1}{x^2}=-\dfrac{kq_2}{(0.04-x)^2},""q_1(0.04-x)^2=-q_2x^2,""\dfrac{0.04-x}{x}=\sqrt{\dfrac{-q_2}{q_1}},""0.04=x\sqrt{\dfrac{-q_2}{q_1}}+x,""x=\dfrac{0.04\ m}{\sqrt{\dfrac{-q_2}{q_1}}+1}=\dfrac{0.04\ m}{\sqrt{\dfrac{-(-6\cdot10^{-9}\ C)}{7\cdot10^{-9}\ C}}+1}=0.021\ m."